OpenStudy (anonymous):

If the height of a refrigerator is 6 feet and the base is 3 feet, find the angle of elevation.

OpenStudy (anonymous):

\(tan\theta = {opposite \over adjacent}\)\[\implies tan\theta = \frac{6}{3} = 2\]\[\implies \theta = tan^{-1}(2)\]

OpenStudy (amistre64):

angle of elevation from which point of view?

OpenStudy (amistre64):

tan is right; since you dont know the hypotenuse right off hand

OpenStudy (anonymous):

Isn't that the length of the fridge though, i.e. 6ft

OpenStudy (amistre64):

6 tall; 3 wide ... got know idea what you need for an elevation tho ....

OpenStudy (anonymous):

I assumed that it was propped against a wall or something. Now I re-read it I'm not got a clue about what I'm answering.

OpenStudy (anonymous):

67.43º 63.43º 26.56º

OpenStudy (anonymous):

it is one of those

OpenStudy (amistre64):

"it" may very well be .. but what is the "it" that we need to look for :)

OpenStudy (amistre64):

tan-1(6/3) gives an answer ... but hard to tell if its right

OpenStudy (anonymous):

In degrees, the arctan of 2 is 63.434949 degrees.